# Stepsize anything: A unified learning rate schedule for budgeted-iteration training

### v1

@@ E.3 Implementation Details
CIFAR10/100 are simple benchmark datasets that are widely used for quick and efficient evaluation of deep learning tasks. CIFAR10 consists of a training split of 50000 examples and a test split of 10000 examples. Each example is a 32$\times$32 color image in 10 classes (airplane, automobile, bird, cat, deer, dog, frog, horse, ship, truck). CIFAR100 comprises 60000 32$\times$32 pixels color images which are divided into 100 classes. Each class contains 600 images and the classes are grouped into 20 super-classes. Each image comes with a ‘fine’ label (the class to which it belongs) and a ‘coarse’ label (the super-class to which it belongs). We use the torch.optim.SGD and torch.optim.AdamW API to configure the optimizer. For vision tasks in Section 4.1, we all adopt the SGD optimizer. For ablation studies in Section 4.3, we adopt the AdamW optimizer. We initialize the learning rate to 1e-1, set the batch size to 128. Each epoch consists of 50000/128 + 1 = 391 steps. We leave the weight decay value as 5e-4, and complete the training task on PyTorch 2.4.1+cu118 with RTX 5880 and RTX 3090.
@@ E.4 Overall experiment results for Vision Classification Tasks
Overall results for the vision classification tasks are presented in Table 7, depicting validation accuracy on vision benchmarks (e.g., CIFAR10/100 and ImageNet) using different architectures (VGG16, ResNet18, ResNet34, ResNet50). It is obvious that our proposed UBA schedule outperform baselines, achieving the highest validation accuracy over both small-scale and large-scale benchmarks across all training budgets.
@@ E.5 Overall experiment results for language models
From Table 8, UBA achieves state-of-the-art performance on approximately 50% of the benchmarks across all scales while the second-best schedule varies. Furthermore, it achieves consistent superior average performance among all baselines. It highlights the stable superiority and reliability of UBA, providing a default-strong choice. Moreover, it demonstrates significant improvements in SciQ-73M($+1.7$) and ARC-E-300M($+2.63$), highlighting its ability to enhance generalization across diverse benchmarks. Besides, in Figure 2, UBA consistently achieves lower training loss and validation loss throughout the training, indicating the efficient training ability and downstream performance enhancement. Notably, while task-specific tuning of $\varphi$ can further improve model performance, we intentionally fix its value across all tasks and architectures to ensure fair evaluation. Remarkably, even with this universal $\varphi$ setting, UBA consistently outperforms baselines on diverse benchmarks, demonstrating inherent robustness to varying benchmarks and model scales.
@@ E.7 Parameter analysis of $\varphi$
E.8 Performance across different periods
@@ Appendix F Related work
Learning rate schedules have been extensively studied to improve training efficiency and model performance. The common scheme is the step decay schedule. Its schedule function can be represented as $g(t)=d_{\lfloor t/t_{s}\rfloor}$, where $t_{s}$ is the predefined decaying step and $d_{i}(d_{i}\geq d_{i+1})$ is predefined decaying scalar. A typical instance decreases the learning rate by a decaying scalar 0.1 after $50\%$ epochs and by a decaying scalar 0.01 after $75\%$ epochs [20]. Then Loshchilov et al. [32] observe that sharp decreases may prevent models from escaping local minima. Thus they propose cosine schedule function as $\frac{1}{2}(1+\cos(\frac{t\pi}{t_{s}}))$, where $t_{s}$ is the predefined stage to restart the learning rate schedule, which has proven effective in transformer training. In addition, cosine schedule is a commonly-used schedule nowadays. Pan et al. [35, 36] further introduce a learning rate schedule called elastic step decay as $\eta_{t}=\eta_{0}/2^{k}$, if $t\in[(1-r^{k})T,(1-r^{k+1})T)$ to accelerate the pre-training for Berts, where $r$ and $k$ are hyper-parameters that control learning rate interval. Hu et al. [24] propose a Warmup-Stable-Decay(WSD) for large model pretraining and promoting continuous training. WSD consists of three consecutive phases: warmup, stable constant, and decay. Luo et al. [33] discover a new schedule that outperforms the cosine schedule, resembling WSD but with lower final loss.

### v4

@@ E.3 Implementation Details
CIFAR10/100 are simple benchmark datasets that are widely used for quick and efficient evaluation of deep learning tasks. CIFAR10 consists of a training split of 50000 examples and a test split of 10000 examples. Each example is a $\times32$ color image in 10 classes (airplane, automobile, bird, cat, deer, dog, frog, horse, ship, truck). CIFAR100 comprises 60000 $\times32$ pixels color images which are divided into 100 classes. Each class contains 600 images and the classes are grouped into 20 super-classes. Each image comes with a ‘fine’ label (the class to which it belongs) and a ‘coarse’ label (the super-class to which it belongs). We use the torch.optim.SGD and torch.optim.AdamW API to configure the optimizer. For vision tasks in Section 4.1, we all adopt the SGD optimizer. For ablation studies in Section 4.3, we adopt the AdamW optimizer. We initialize the learning rate to 1e-1, set the batch size to 128. Each epoch consists of 50000/128 + 1 = 391 steps. We leave the weight decay value as 5e-4, and complete the training task on PyTorch 2.4.1+cu118 with RTX 5880 and RTX 3090.
@@
Our implementation strictly follows the PyTorch official API, i.e. it is inherited from the base class
Ψtorch.optim.lr_scheduler.LRScheduler
. It maintains full compatibility with all
Ω
torch.optim
optimizers. The complete production-ready code is publicly available.
@@ E.4 Overall experiment results for Vision Classification Tasks
Overall results for the vision classification tasks are presented in Table 7, depicting validation accuracy on vision benchmarks (e.g., CIFAR10/100 and ImageNet) using different architectures ( ResNet18, ResNet34, ResNet50). We conduct several independent runs for error bars. Due to computational constraints, we use two independent experimental results for ImageNet-ResNet50. It is obvious that our proposed UBA schedule outperform baselines, achieving the highest validation accuracy over both small-scale and large-scale benchmarks across all training budgets.
@@ E.5 Overall experiment results for language models
From Table 8, UBA achieves state-of-the-art performance on approximately 50% of the benchmarks across all scales while the second-best schedule varies. Furthermore, it achieves consistent superior average performance among all baselines. It highlights the stable superiority and reliability of UBA, providing a default-strong choice. Moreover, it demonstrates significant improvements in SciQ-73M(+1.7+1.7) and ARC-E-300M(+2.63+2.63), highlighting its ability to enhance generalization across diverse benchmarks. Besides, in Figure 2, UBA consistently achieves lower training loss and validation loss throughout the training, indicating the efficient training ability and downstream performance enhancement. Notably, while task-specific tuning of $\varphi$ can further improve model performance, we intentionally fix its value across all tasks and architectures to ensure fair evaluation. Remarkably, even with this universal $\varphi$ setting, UBA consistently outperforms baselines on diverse benchmarks, demonstrating inherent robustness to varying benchmarks and model scales.
@@ E.7 Parameter analysis of $\varphi$
E.8 Performance across different initial learning rates
Since each schedule will have a different optimal learning rate in practice, to thoroughly address this concern, we conducted additional experiments to evaluate the performance of all schedules with different learning rates.
We conducted auxiliary experiments on CIFAR100-ResNet34 for 3 runs. For computational efficiency, these experiments were run for slightly fewer epochs than the main experiments, with results reported over three independent runs. Through current observation and search, we identified 0.1 as the proper initial learning rate for most schedules. We use 0.01 and 0.001 to confirm this choice maximizes validation accuracy across methods. Additional smaller or larger values such as 0.3 or 0.0001 were explored in preliminary studies, showing consistent trends but lower peak accuracy. Thus we do not choose them for comparison.
The experimental results are listed in the following table 11. We report results of different base learning rate and compare the performance with optimal learning rate of each schedule.
Table 11: Performance across different initial learning rates.
Dataset Network Schedule Validation Accuracy
lr=0.1 lr=0.01 lr=0.001
[…]
E.9 Performance across different periods
@@ Appendix F Related work
Learning rate schedules have been extensively studied to improve training efficiency and model performance. The common scheme is the step decay schedule. Its schedule function can be represented as $g(t)=d_{\lfloor t/t_{s}\rfloor}$, where $t_{s}$ is the predefined decaying step and $d_{i}(d_{i}\geq d_{i+1})$ is predefined decaying scalar. A typical instance decreases the learning rate by a decaying scalar 0.1 after $50\%$ epochs and by a decaying scalar 0.01 after $75\%$ epochs [21]. Then Loshchilov et al. [33] observe that sharp decreases may prevent models from escaping local minima. Thus they propose cosine schedule function as $\frac{1}{2}(1+\cos(\frac{t\pi}{t_{s}}))$, where $t_{s}$ is the predefined stage to restart the learning rate schedule, which has proven effective in transformer training. In addition, cosine schedule is a commonly-used schedule nowadays.
Pan et al. [36] propose Eigencurve, and demonstrate that minimax-optimal convergence can be achieved for quadratic objectives when the Hessian’s eigenvalue distribution exhibits substantial skewness. Pan et al. [37] further introduce a learning rate schedule called elastic step decay as $\eta_{t}=\eta_{0}/2^{k}$, if $t\in[(1-r^{k})T,(1-r^{k+1})T)$ to accelerate the pre-training for Berts, where $r$ and $k$ are hyper-parameters that control learning rate interval. Defazio et al. [11] developed the theoretically-grounded framework for learning rate scheduling, establishing optimal linear decay schedules and refinements through rigorous mathematical derivation. Their work additionally provides the most extensive empirical evaluation of scheduling approaches to date. Hu et al. [25] propose a Warmup-Stable-Decay(WSD) for large model pretraining and promoting continuous training. WSD consists of three consecutive phases: warmup, stable constant, and decay. Luo et al. [34] discover a new schedule that outperforms the cosine schedule, resembling WSD but with lower final loss.